# Mellivora Capensis: A Backdoor-Free Training Framework on the Poisoned Dataset without Auxiliary Data

### v2

@@ 4.1 Overview of MeCa
4.2 Detailed Design of MeCa
@@
4.2.1 Exploring the Relationship between Perturbation and Backdoor
@@
Our investigation begins with a review of the backdoor attack process, which is similar with [50]. Let $\mathcal{D}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$ represent a clean training set, and let $C:\mathcal{X}\rightarrow\mathcal{Y}$ denote the functionality of the target neural network. Each image $\boldsymbol{x}_{i}$ in $\mathcal{D}$ satisfies $\boldsymbol{x}_{i}\in\mathcal{X}=[0,1]^{C\times W\times H}$, and its corresponding label $y_{i}\in\mathcal{Y}=\{1,\ldots,K\}$, where $K$ is the total number of classes. To initiate an attack, adversaries poison a selection of clean samples $\mathcal{D}_{p}$ using a covert transformation $T(\cdot)$. These poisoned samples are then combined with the clean dataset prior to training the compromised model. This process can be formalized as $\mathcal{D}_{t}=\mathcal{D}\cup\mathcal{D}_{p}$, where $\mathcal{D}_{p}=\{(x_{i}^{\prime},y_{t})|x_{i}^{\prime}=T(x_{i}),(x_{i},y_{i})\in\mathcal{D}_{p}\}$. Various methods have been devised to make $T(\cdot)$ more subtle and harder to detect, or to reduce the assumptions about the adversary’s capabilities, such as the poison ratio $pr=|\mathcal{D}_{p}|/|\mathcal{D}|$. Despite these variations, the ultimate goal remains the same [50]: manipulating the neural network by training a malicious model:
@@
where $N_{b}=|\mathcal{D}|$ and $N_{p}=|\mathcal{D}_{p}|$.
@@
$\mathcal{L}_{kl}(f(x),f(x\times(1-\hat{m})+\hat{m}\times\delta))).$ (3)
@@
To further explain this intriguing phenomenon, we propose the corresponding proof based on the recent studies [50] to analyze the characteristics of poisoned samples. We start with the regression solution for the Neural Tangent Kernel (NTK) model as presented in [53, 50, 54]:
@@
Consider a specific backdoor sample $\boldsymbol{x}^{\prime}$, which is crafted as a mixture of a clean sample $\boldsymbol{x}$ and a target pattern $\boldsymbol{t}$, i.e., $\boldsymbol{x}^{\prime}=(1-m)\odot\boldsymbol{x}+m\odot\boldsymbol{t}$, where $\odot$ denotes element-wise multiplication, and $m$ controls the blending ratio between $\boldsymbol{x}$ and $\boldsymbol{t}$.
When $N_{p}$ approaches $N_{b}$ (i.e., the poisoning ratio is close to 50%), the attack’s efficacy is maximized. In this case, the expression is:
@@
To get the inequality blow:
$\phi_{t}(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime})\geq\frac{1}{2}$ (8)
We need to have:
$\sum_{i=1}^{N_{b}}Q\left(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime},\boldsymbol{x}_{\boldsymbol{i}}\right)\leq\sum_{i=1}^{N_{p}}Q\left(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime},\boldsymbol{x}_{\boldsymbol{i}}^{\prime}\right)$ (9)
@@
Figure 5: Loss landscape of clean sample and poisoned sample.
Theorem 1. Let $\mathcal{D}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N_{b}}$ be a clean dataset consisting of $N_{b}$ samples, and $\mathcal{D}_{p}=\{(\boldsymbol{x}_{i}^{\prime},y_{t}^{\prime})\}_{i=1}^{N_{p}}$ be a poisoned dataset containing $N_{p}$ samples, both independently and identically distributed (i.i.d.) according to a uniform distribution across $K$ classes. Suppose a deep neural network $f(\cdot;\theta)$ is formulated as a multivariate kernel regression using a Radial Basis Function (RBF) kernel, and shares the same objective as the attackers. For a given attacked sample $\boldsymbol{x}^{\prime}=(\mathbf{1}-\boldsymbol{m})\times\boldsymbol{x}+\boldsymbol{m}\times\boldsymbol{t}$, where $\mathbf{1}$ is a vector of ones, $\boldsymbol{m}$ is a binary mask, $\boldsymbol{x}$ is the original sample, and $\boldsymbol{t}$ is the trigger pattern, we have: $\lim_{N_{p}\to N_{b}}C(\boldsymbol{x}^{\prime}+\delta)=y_{t}$, where $C(\cdot)$ denotes the classification function of the network, $\delta$ represents a perturbation added to the attacked sample, and $y_{t}$ is the target class label.
The theorem above demonstrates that when the size of $\mathcal{D}_{p}$ approaches that of the clean dataset, poisoned samples exhibit robustness against potential perturbations. To gain an intuitive understanding, we visualize the loss landscapes of both clean and poisoned samples (using the Blend attack [25]), as illustrated in Fig. 5. Notably, the loss of a clean sample shows steeper changes when perturbations are applied in the direction of the gradient. This observation further corroborates the previous theoretical analysis.
@@
1:  $Q=[\ ]$ // Initialize empty list $Q$
@@
6:     $\hat{x}=x\times(1-\hat{m})+\hat{m}\times\delta$
7:     // Calculate $KL$ divergence of each sample
8:     $Q.append(\mathcal{L}_{kl}(f(x),f(\hat{x})))$
@@
Figure 6: ASR and ACC of the model during backdoor enhance training on CIFAR10 with ResNet18.
4.2.2 Dataset Partition
Based on the findings in the above section, we plan to partition the untrusted dataset roughly by using the characteristic that the poisoned samples perform greater robustness against potential perturbations than the clean samples. However, as shown in Fig.4, these poisoned samples are still mixed with some clean ones. Therefore, in this section, we just pick out a small number of clean samples iteratively during the training stage rather than precisely selecting all the poisoned samples. Note that we do not require any auxiliary clean dataset.
@@ 4.2.2 Dataset Partition
4.2.3 Backdoor Enhancement and Standard Training
Note that a small number of clean samples $\mathcal{D}_{c}$ means nothing to train a clean model. However, motivated by [18], a small number of clean samples is enough to unlearn the functionality of a model on the main task. That’s to say, we can obtain a model $\hat{f}_{\theta}$ with a high attack success rate but low accuracy by unlearning some clean samples. To achieve a higher ASR and lower ACC, we train a backdoor-enhanced model iteratively by unlearning a small number of clean samples and learning the rest poisoned dataset. Concretely, for each iteration, $p$ of samples with the lowest consistency will be labeled as clean samples $\mathcal{D}_{c}$ and unlearned:
@@ 4.2.3 Backdoor Enhancement and Standard Training
During this process, more clean samples will be picked out for unlearning, and the model will be trained on a dataset with an increasing poison ratio. As illustrated in Fig.6, ASR and ACC of the model continue to increase and decline, respectively. Algorithm 2 shows the iterative backdoor-enhanced training process.
@@ 5 Experiments
5.1 Experimental Settings
5.1.1 Datatsets
@@ 5.1.3 Experimental Environment
5.1.4 Backdoor Attacks
We employ $8$ state-of-the-art backdoor attacks to evaluate the proposed MeCa in our experiments, including $3$ dirty-label attacks: BadNet [12], Blend [25], and PhysicalBA [46], one clean-label attack: SIG [48], one input-specific-trigger attack: WaNet [47], and $3$ adaptive attacks: TaCT [43], AdaptiveBlend [49] and AdaptivePatch [49]. The detailed configurations of these backdoor attacks are shown in Table I.
BadNet 0.05 3*3 white square Static Dirty No / 1 /
Blend 0.05 HelloKity , alpha=0.2 Static Dirty No / 1 /
SIG 0.05 Sinusoidal signal Static Clean No / 1 /
WaNet 0.05 Warping-based triggers Dynamic Dirty No / 1 0.05
PhysicalBA 0.05 Firefox Static Dirty No / 1 /
TaCT 0.05 3*3 white square Static Dirty Yes 0 1 0.01
AdaptiveBlend 0.05 HelloKity , alpha=0.2 Static Dirty Yes / 1 0.01
AdaptivePatch 0.05 Firefox*4 Static Dirty Yes / 1 0.01
5.1.5 Backdoor Defense
@@
  • DBR: DBR tries to pick out with high sensitivity to “rotate” and “affine” transformation. We set clean ratio and poison ratio to $0.4$ and $0.05$, respectively, and unlearn the poisoned samples for $5$ epochs with $0.0001$ unlearning rate.
@@
5.1.6 Evaluation Metrics
We employ the attack success rate (ASR) and accuracy (ACC) to evaluate the performance of the proposed backdoor defense framework. ASR indicates the ratio of the triggered samples that are misclassified as the target label. ACC represents the accuracy of the model on the clean samples. The lower ASR and the higher ACC indicate the better performance of the backdoor defense mechanism.
@@ 5.1.6 Evaluation Metrics
TABLE II: The comparison results (%) between our MeCa and $7$ state-of-the-art backdoor defenses methods against $5$ state-of-the-art backdoor attacks on ResNet18 with $5\%$ poisoned ratio.
@@
5.2 Comparative Evaluation of MeCa and Existing Defenses
@@
(3) The existing defense methods perform unsteadily for the same backdoor attack on the different datasets. For example, for DBR under the SIG, the ASR on four datasets is $0.28\%$, $93.52\%$, $80.10\%$, and $0.00\%$, respectively. For ABL under the Blend, the ASR on four datasets is $79.22\%$, $5.71\%$, $1.21\%$, and $0.00\%$, respectively. Since the complexity, scale, and class numbers of the different datasets vary, the convergence speed of attack is also different. This results in a high false positive rate for poisoned sample identification when the backdoor task has not been learned. In this case, some clean samples may be picked out. Unlearning these clean samples will significantly influence the model performance.
(4) All the existing defense methods have an unstable performance against different backdoor attacks. They perform satisfactorily on some backdoor attacks and may be ineffective against the other backdoor attacks. This greatly reduces the availability of these defense methods in real-world scenarios because it is impractical for the defender to know the backdoor methods in advance. For example, ABL performs well on the BadNet attack but has an unsatisfactory performance on the SIG attack. Note that we have observed the different performance of these defense methods in different papers because their hyperparameters are different. We argue that it’s not practical to set different hyperparameters for different attacks because defenders have no knowledge about the potential attacks. Therefore, we use the same hyperparameter configurations for other defenses as well as MeCa.
(5) Compared with the $7$ state-of-the-art backdoor defense techniques, our MeCa can achieve the lowest ASR (the highest ASR is just $2.64\%$) while maintaining a satisfactory performance on the clean samples. Moreover, the proposed MeCa has a more stable performance on different datasets and various backdoor attacks.
TABLE III: The comparison results (%) between our MeCa and $7$ state-of-the-art backdoor defenses methods against $3$ state-of-the-art adaptive attacks on ResNet18 with $5\%$ poisoned ratio.
@@
5.3 The Resistance to Adaptive Attacks

### v3

@@ IV-A Overview of MeCa
IV-B Detailed Design of MeCa
@@
IV-B1 Exploring the Relationship between Perturbation and Backdoor
@@
Our investigation begins with a review of the backdoor attack process, which is similar with [60]. Let $\mathcal{D}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$ represent a clean training set, and let $f\in\mathcal{F}:\mathcal{X}\rightarrow\mathcal{Y}$ denotes the functionality of the target neural network. Each image $\boldsymbol{x}_{i}$ in $\mathcal{D}$ satisfies $\boldsymbol{x}_{i}\in\mathcal{X}=[0,1]^{C\times W\times H}$, and its corresponding label $y_{i}\in\mathcal{Y}=\{1,\ldots,K\}$, where $K$ is the total number of classes. To initiate an attack, adversaries poison a selection of clean samples $\mathcal{D}_{p}$ using a covert transformation $T(\cdot)$. These poisoned samples are then combined with the clean dataset prior to training the compromised model. This process can be formalized as $\mathcal{D}_{t}=\mathcal{D}\cup\mathcal{D}_{p}$, where $\mathcal{D}_{p}=\{(x_{i}^{\prime},y_{t})|x_{i}^{\prime}=T(x_{i}),(x_{i},y_{i})\in\mathcal{D}_{p}\}$. Various methods have been devised to make $T(\cdot)$ more subtle and harder to detect, or to reduce the assumptions about the adversary’s capabilities, such as the poison ratio $pr=|\mathcal{D}_{p}|/|\mathcal{D}|$. Despite these variations, the ultimate goal remains the same [60]: manipulating the neural network by training a malicious model:
@@
where $N_{b}=|\mathcal{D}|$ and $N_{p}=|\mathcal{D}_{p}|$ and $\mathcal{L}$ typically stands for cross entropy.
@@
$\mathcal{L}_{kl}(f(x),f(x\odot(1-\hat{m})+\hat{m}\odot\delta))).$ (3)
@@
To further explain this intriguing phenomenon, we propose the corresponding proof based on the recent studies [60] to analyze the characteristics of poisoned samples. We start with the regression solution for the Neural Tangent Kernel (NTK) model. The NTK is a theoretical framework that describes the training dynamics of deep neural networks in the limit of infinite width [63]. It establishes that for an infinitely wide network trained with gradient descent, the learning process simplifies to a kernel regression problem where the kernel remains constant throughout training. The NTK can also be used to formally analyze the characteristics of poisoned samples as presented in [64, 60, 65]
@@
Consider a specific backdoor sample $\boldsymbol{x}^{\prime}$, which is crafted as a mixture of a clean sample $\boldsymbol{x}$ and a target pattern $\boldsymbol{t}$. When $N_{p}$ approaches $N_{b}$ (i.e., the poison ratio is close to 50%), the attack’s efficacy is maximized. In this case, the expression is:
@@
To satisfy the inequality below:
$\phi_{t}(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime})\geq\frac{1}{2},$ (8)
a sufficient condition is that the sum of kernels for clean samples is less than or equal to that of poisoned samples:
$\sum_{i=1}^{N_{b}}Q\left(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime},\boldsymbol{x}_{\boldsymbol{i}}\right)\leq\sum_{i=1}^{N_{p}}Q\left(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime},\boldsymbol{x}_{\boldsymbol{i}}^{\prime}\right).$ (9)
@@
Figure 5: Loss landscapes for a representative clean and poisoned sample.
Theorem 1. Let $\mathcal{D}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N_{b}}$ be a clean dataset consisting of $N_{b}$ samples, and $\mathcal{D}_{p}=\{(\boldsymbol{x}_{i}^{\prime},y_{t}^{\prime})\}_{i=1}^{N_{p}}$ be a poisoned dataset containing $N_{p}$ samples, both independently and identically distributed (i.i.d.) according to a uniform distribution across $K$ classes. Suppose a deep neural network $f(\cdot;\theta)$ is formulated as a multivariate kernel regression using a Radial Basis Function (RBF) kernel, and shares the same objective as the attackers. For a given attacked sample $\boldsymbol{x}^{\prime}=(\mathbf{1}-\boldsymbol{m})\odot\boldsymbol{x}+\boldsymbol{m}\odot\boldsymbol{t}$, where $\mathbf{1}$ is a vector of ones, $\boldsymbol{m}$ is a binary mask, $\boldsymbol{x}$ is the original sample, and $\boldsymbol{t}$ is the trigger pattern, we have: $\lim_{N_{p}\to N_{b}}C(\boldsymbol{x}^{\prime}+\delta)=y_{t}$, where $C(\cdot)$ denotes the classification function of the network, $\delta$ represents a perturbation added to the attacked sample, and $y_{t}$ is the target class label.
The theorem above demonstrates that when the size of $\mathcal{D}_{p}$ approaches that of the clean dataset, poisoned samples exhibit robustness against potential perturbations. Note that we must acknowledge there is an apparent gap between the theoretical assumption in Theorem 1, which considers an asymptotic case where $N_{p}\rightarrow N_{b}$ (poison ratio approaches 50%), and the lower, more practical ratios in our experiments. Our intention with the theorem was to provide a theoretical intuition in an analytically tractable setting to explain why backdoor features become exceptionally robust. The theorem demonstrates that under these conditions, the backdoor becomes the dominant feature. Our extensive empirical results then demonstrate that this principle of differential robustness holds true and is practically effective even at much lower poisoning rates.
To gain a further intuitive understanding, we visualize the loss landscapes of both clean and poisoned samples (using the Blend attack [32]), as illustrated in Fig. 5. Notably, the loss of a clean sample shows steeper changes when perturbations are applied in the direction of the gradient. This observation further corroborates the previous theoretical analysis [66] in which the author found that the square sum of the eigenvalues of the Hessian matrix of the benign model is higher than that of the backdoored one. To further enlarge such a sensitivity (robustness) gap, we apply the worst-case perturbations (adversarial perturbation) to make poisoned samples as distinguishable as possible.
@@
1:Q=[]Q=[\ ] // Initialize empty list $Q$
@@
6:  $\hat{x}=x\odot(1-\hat{m})+\hat{m}\odot\delta$
7:  // Calculate KLKL divergence of each sample
[…]
@@
Figure 6: ASR and ACC of the model during backdoor-enhanced training on CIFAR10 with ResNet18.
IV-B2 Dataset Partition
Based on the findings in the above section, we plan to partition the untrusted dataset roughly by using the characteristic that the poisoned samples perform greater robustness against potential perturbations than the clean samples. However, as shown in Fig.4, these two histograms (y-axis means number and x-axis means KL divergence) illustrate that poisoned samples are still mixed with some clean ones. Therefore, in this section, we only pick out a small number of clean samples iteratively during the training stage rather than precisely selecting all the poisoned samples.
@@ IV-B2 Dataset Partition
IV-B3 Backdoor Enhancement and Standard Training
Note that a small number of clean samples $\mathcal{D}_{c}$ means nothing to train a clean model. However, motivated by [25], a small number of poisoned samples is enough to unlearn the functionality of the backdoor. We instead unlearn the main task on the “possibly” backdoored model with selected benign samples. That’s to say, we can obtain a model $\hat{f}_{\theta}$ with a high ASR but low ACC by unlearning some clean samples, which can be recognized as a poisoned sample detector. To achieve a higher ASR and lower ACC, we train a backdoor-enhanced model iteratively by unlearning a small number of clean samples and learning the rest poisoned dataset. Concretely, for each iteration, $p$ of samples with the lowest consistency will be labeled as clean samples $\mathcal{D}_{c}$ and unlearned:
@@ IV-B3 Backdoor Enhancement and Standard Training
During this process, more and more clean samples will be picked out for unlearning with the decrease of $p$, and the model will also be trained on a dataset with an increasing poison ratio. The rationale for designing for adopting such a dynamic process is that the $p$ must be small for its weak ability to detect poisoned samples. While unlearning with more benign samples and learning with more poisoned samples, the ability of the detection increases as well and a higher $p$ is set to pick out more benign samples. As illustrated in Fig.6, ASR and ACC of the model continue to increase and decline, respectively. Algorithm 2 shows the iterative backdoor-enhanced training process.
@@
BadNet 5% 3*3 white square Static Dirty No / 1 /
Blend 5% HelloKity , alpha=0.2 Static Dirty No / 1 /
SIG 5% Sinusoidal signal Static Clean No / 1 /
WaNet 5% Warping-based triggers Dynamic Dirty No / 1 5%
PhysicalBA 5% Firefox Static Dirty No / 1 /
TaCT 5% 3*3 white square Static Dirty Yes 0 1 0.01
AdaptiveBlend 5% HelloKity , alpha=0.2 Static Dirty Yes / 1 0.01
AdaptivePatch 5% Firefox*4 Static Dirty Yes / 1 0.01
@@ V Experiments
V-A Experimental Settings
V-A1 Datatsets
@@ V-A3 Experimental Environment
V-A4 Backdoor Attacks
We employ $8$ state-of-the-art backdoor attacks to evaluate the proposed MeCa in our experiments, including $3$ dirty-label attacks: BadNet [13], Blend [32], and PhysicalBA [56], one clean-label attack: SIG [58], one input-specific-trigger attack: WaNet [57], and $3$ adaptive attacks: TaCT [53], AdaptiveBlend [59] and AdaptivePatch [59]. The poison ratio is set to $5\%$, aligning with most previous works [25, 53, 26, 70, 71, 31]. The detailed configurations of these backdoor attacks are shown in Table I.
V-A5 Backdoor Defense
@@
  • DBR: DBR tries to pick out with high sensitivity to “rotate” and “affine” transformation. We set clean ratio and poison ratio to $40\%$ and $5\%$, respectively, and unlearn the poisoned samples for $5$ epochs with $0.0001$ unlearning rate.
@@
  • MeCa: For our defense, we only apply a single-step backward to obtain the perturbation bounded by $0.001$, and the patch size is set to 2 for all attacks. Note that both of the constraints make it impossible to affect the functionality of the backdoor while slightly shifting the logits distribution of the clean samples. The unlearning coefficient is set to 0.05 across any datasets.
V-A6 Evaluation Metrics
We employ the ASR and ACC to evaluate the performance of the proposed backdoor defense framework. ASR indicates the ratio of the triggered samples that are misclassified as the target label. ACC represents the accuracy of the model on the clean samples. The lower ASR and the higher ACC indicate the better performance of the backdoor defense mechanism.
@@ V-A6 Evaluation Metrics
TABLE II: The comparison results (%) between our MeCa and 77 state-of-the-art backdoor defenses methods against $5$ state-of-the-art backdoor attacks on ResNet18 with $5\%$ poisoned ratio.
@@
V-B Comparative Evaluation of MeCa and Existing Defenses
@@
(3) The existing defense methods perform unsteadily for the same backdoor attack on the different datasets. For example, for DBR under the SIG, the ASR on four datasets is $0.28\%$, $93.52\%$, $80.10\%$, and $0.00\%$, respectively. For ABL under the Blend, the ASR on four datasets is $79.22\%$, $5.71\%$, $1.21\%$, and $0.00\%$, respectively. Since the complexity, scale, and class numbers of the different datasets vary, the convergence speed of the attack is also different. This results in a high false positive rate for poisoned sample identification when the backdoor task has not been learned. In this case, some clean samples may be picked out. Unlearning these clean samples will significantly influence the model’s performance.
(4) All the existing defense methods have an unstable performance against different backdoor attacks. They perform satisfactorily on some backdoor attacks and may be ineffective against other backdoor attacks. This greatly reduces the availability of these defense methods in real-world scenarios because it is impractical for the defender to know the backdoor methods in advance. For example, ABL performs well on the BadNet attack but has an unsatisfactory performance on the SIG attack. Note that we have observed the different performance of these defense methods in different papers because their hyperparameters are different. We argue that it’s not practical to set different hyperparameters for different attacks because defenders do not know the potential attacks. Therefore, we use the same hyperparameter configurations for different strategies, including MeCa.
(5) Compared with the $7$ state-of-the-art backdoor defense techniques, our MeCa can achieve the lowest ASR (the highest ASR is just $2.64\%$) while maintaining a satisfactory performance on the clean samples. While some defenses reduce ASR, they do so at the cost of rendering the model practically useless, thereby failing to maintain model availability. This can be attributed to their over-unlearning characteristics to obtain low ASR, with a suboptimal trade-off between robustness and utility. Moreover, the proposed MeCa has a more stable performance on different datasets and various backdoor attacks.
TABLE III: The comparison results (%) between our MeCa and 77 state-of-the-art backdoor defenses methods against $3$ state-of-the-art adaptive attacks on ResNet18 with $5\%$ poisoned ratio.
@@
V-C The Resistance to Adaptive Attacks